# Bound-electron self-energy calculations in Feynman and Coulomb gauges: detailed analysis

### v1

@@ III Computational details
The calculations are carried out by means of two approaches: the FBS method within the DKB basis and the GF method. The methods differ only in the treatment of the terms evaluated in coordinate space. The momentum-space calculations are the same in both approaches.
@@ III.1 Momentum-space calculations
The contributions evaluated in momentum space include the zero- and one-potential terms resulting from the renormalization procedure, as well as the quasi-two-potential terms corresponding to the SC convergence-acceleration scheme. All the integrals in the zero-, one-, and quasi-two-potential terms are calculated using the Gauss-Legendre quadratures. Compared to the Feynman gauge, the zero-potential contribution in the Coulomb gauge contains an additional integration over the Feynman parameter $x$, see Eqs. (17), (56), (57). The coefficient $F_{2}$ in Eq. (57) must be calculated as a principal-value integral. From a numerical point of view, it is necessary to isolate the singularity. We do that by adding and subtracting the same integrand but with the numerator evaluated at $x=a$:
@@
The first term in the right-hand side of Eq. (41) becomes regular. To avoid numerical problems at $x=a$, the exact expression can be replaced with its Taylor series in a small vicinity of this point. The second term in Eq. (41) is calculated analytically. The domain of integration over $p$ is divided into two parts, $p\leq p_{0}$ and $p\geq p_{0}$, where $p_{0}=\alpha Z$. The integration over the second domain can be carried out using a change of variables of the form $t=\alpha Z/p$, as a result of which the ray $\alpha Z,\infty)$, is mapped onto the interval $[1,0]$. Similar methods can be used to calculate $\Delta\tilde{\varepsilon}_{a}^{2\text{p}}$ in momentum space. To avoid possible numerical problems due to large cancellations, one should use the Taylor-series expansion for the corresponding integrands.
The evaluation of the one-potential contributions involves the integrations over the variables $p$, $p^{\prime}$, and $z$ and a number of Feynman parameters. When integrating over the Feynman parameters in both gauges, to avoid possible numerical errors, the Taylor series for the integrands are employed. In the Coulomb-gauge case we use the results of Ref. OP_CG, but expand all the coefficients $F_{1}$ - $F_{21}$ at ‘‘suspicious’’ points. The expression (20) has an integrable singularity in $q=0$ at $p=p^{\prime}$ and $z=1$. Therefore it is necessary to properly choose integration nodes for the variable $z$ to ensure the integral convergence. For this goal, we use the transformation $z=1-2t^{2}$. To calculate the integrals over the variables $p$ and $p^{\prime}$, we change the variables according to $P=(p+p^{\prime})/\sqrt{2}$ and $P^{\prime}=(p-p^{\prime})/\sqrt{2}$. The integral over $P$ is calculated similar to the integral over $p$ for the zero-potential contribution, but the separation point $p_{0}=2\alpha Z$ is chosen instead. The integral over $P^{\prime}$ is taken within the finite interval [P,P][-P,P].
To estimate an uncertainty of the momentum-space calculations, the number of quadrature nodes and position of the separation points $p_{0}$ is varied.
@@ III.2 Coordinate-space calculations
The evaluation of all coordinate-space contributions, many-potential, two-potential, and (quasi-) three-plus-potential, has many common points. Let us illustrate them with an example of the many-potential contribution. Its evaluation essentially comes down to the calculation of the three integrals: over the radial variables $r_{1}$ and $r_{2}$, see Eqs. (71) and (72), and over the energy parameter $\omega$. Taking into account the exponential decrease of bound-electron wave functions at large distances, all radial integrations are performed over the finite interval $0,R_{max}]$. The value of the parameter $R_{max}$ is chosen so that the results are independent of it within the desired accuracy. All the radial integrations are carried out using the Gauss–Legendre quadratures. However, the specific details differ slightly in the DKB and GF approaches.
In the DKB approach, the domain $0,R_{max}]$ is divided into a finite number of intervals by the nodes used to construct the B splines. The integration method for the double integral over $r_{1}$ and $r_{2}$ depends on whether these variables belong to the same interval or not. If not, the integrations are carried out independently using the same quadrature for both variables. However, when $r_{1}$ and $r_{2}$ fall into the same interval, a more delicate integration is employed to take into account the fact that the integrand depends on $r_{<}=\min\{r_{1},r_{2}\}$ and $r_{>}=\max\{r_{1},r_{2}\}$. Namely, for some fixed value of $r_{1}$ the corresponding interval is divided into two subitervals lying to the left and to the right from $r_{1}$. The integration over $r_{2}$ is performed by using the Gauss-Legendre quadratures for both subintervals. Note that within the DKB approach, integration over radial variables turns out to be ‘‘tied’’ to the size of the basis $N$. As the basis-set size grows, the number of the nodes also increases, that leads to the improvement of the integration accuracy. Thus, the single parameter $N$ affects both the accuracy of the Green’s function representation and the integration accuracy.
In the GF approach, the radial integrations become more complicated due to the aforementioned discontinuities of the electron Green’s functions. The integrations over $x$ and $y$ in Eq. ((25)) cannot be performed independently of the integrations over $r_{1}$ and $r_{2}$. To evaluate the radial integrals, we employ a variation of the method described, e.g., in Ref. sym12050800. Namely, a hierarchy of the radial integration grids is prepared, with each subsequent grid being finer and "lying" inside the previous one. Then, all the radial integrations are arranged in the following order: $r_{1}$, $r_{2}$, $x$, and $y$. Therefore, the finest grid corresponds to the variable $y$. The accuracy of the GF approach is determined mainly by the accuracy of the radial integrations.
@@
Figure 4: Deformed contour (blue line) in the complex plane $\omega$. Singularities and branching cuts of the electron Green’s function (black lines and black dots) and photon propagator (black dashed lines) are shown.
The photon and electron propagators have a complicated analytical structure if considered as functions of the energy parameter $\omega$ in the complex plane. They are two-valued functions and are defined on the corresponding Riemann surfaces with two sheets (for each propagator). Fig. 4 shows the complex plane of the integration variable $\omega$, the branch points with the cuts starting from them, and the poles of the electron propagator corresponding to the bound states. In the expression for the SE correction (3), the integration contour goes along the real axis, but it can be bended to ‘‘simplify’’ the calculations. Because of the zero photon mass, the contour is squeezed between two infinitely close branching points. The introduction of a finite mass would allow to push them apart, which can be easily demonstrated in the Feynman gauge. Therefore, the contour is always obliged to pass through $\omega=0$. In principle, there is no other restrictions on the contour. Thus, it can be changed as needed, respecting the analytical structure of the integrand.
In this work, the approach proposed in Ref. base_1999 is used. Namely, we employ the transformation that results in the contour shown in Fig. 4. The original contour along the real axis is rotated and deformed. The rotation is done to avoid fast oscillations of the integrand encountered on the real axis: on the vertical parts of the deformed contour the integrand decays exponentially. The deformation of the contour near zero and its shift to the right from the imaginary axis is done to avoid going around the poles of the electron Green’s function, which correspond to the bound states with binding energies no less than that of the state under consideration. The integral over $\omega$ is evaluated using the Simpson quadratures. The number of integration points is increased until the required accuracy is achieved.
@@ III.3 Extrapolation of the PW-expansion sums
We obtain a coordinate-space contribution as a series in $\kappa$, which should inevitably be truncated at some $|\kappa_{\text{max}}|$. To estimate the remainder of the series, we must resort to an extrapolation procedure. The extrapolation methods discussed below are applied to all treated coordinate-space contributions.
Here and in what follows, when discussing partial sums, we will denote by $k>0$ the sum of terms of the PW-expansion with $|\kappa|=k$. When using the DKB method, the individual contributions strongly depend on the basis-set size $N$, therefore, an extrapolation over this parameter is also required. In this case, the most stable results are obtained by first extrapolating with respect to the basis size. Let us denote as $S_{k}(N)$ the partial sum obtained for the given basis-set size $N$. For each $k$, the extrapolation over $N$ is performed using a non-linear least-squares method using an ansatz of the form:
@@
where $S_{k}$ is the extrapolated to the infinite basis set partial sum, and $A_{k}$ and $B_{k}$ are numerical coefficients. In the GF approach, the sums $S_{k}$ are obtained directly from the calculations, and this extrapolation is not required.
To extrapolate over $k$ in both DKB and GF approaches, we implement the following anzatz for the partial sums $S_{k}$:
@@
where the parameter $S_{\infty}$ and coefficients $C_{i}$ for $i=2\ldots m$ (with $m$ being typically about $5$) are obtained by a statistical method. Namely, we select a random ‘‘kappa’’ subset $\{k_{i}\}_{i=1}^{m}$ a certain number $N_{\text{iter}}=100-1000$ of times. An additional condition is imposed that there are no consecutive $k$ in the chosen subset. We then construct and solve the system of $m$ linear equations of the form
$\tilde{S}(k_{i})=S_{k_{i}},\;i=1\dots m$ (44)
to determine $m$ variables $S_{\infty}$, $C_{2}$, … $C_{m}$. This procedure is repeated $N_{\text{iter}}$ times, resulting in a set of approximations for $S_{\infty}$. Evaluating the mean value $\overline{S}_{\infty}$ and variance $\sigma_{k}$ one obtains the desired extrapolated sum and an estimation for its uncertainty.
Within DKB approach, the uncertainties of extrapolation in k, $\sigma_{k}$, and extrapolation to an infinite basis set, $\sigma_{n}$, are summed quadratically
@@
Example of the use of the two-step extrapolation procedure can be found in the ‘‘results’’ section below.
@@ IV Results
We conduct two sets of SE correction calculations. First we consider the $1S_{1/2}$ state in three hydrogen-like ions: argon ($Z=18$), xenon ($Z=54$), and uranium ($Z=92$). We adopt the results from the work hen as a benchmark for the comparison. For this reason, the values of nuclear radii as well as the values of fundamental constants, e.g. the fine-structure constant $\alpha$ and the Hartree energy in eV, are taken to be the same as there. The obtained data are used to analyze the convergence of the many-potential contribution.
We then reproduce the results from the work se_yerokh, the rms radii and models of the nuclei are chosen to coincide with the corresponding work. In that part, the SE correction is presented in terms of the function $F(\alpha Z)$, defined as:
@@
We use the numerical value for $\alpha$ from the latest CODATA (2018) codata_2018. The data obtained in the second part of the work are used to analyze acceleration-convergence schemes.
@@ IV.1 Many-potential term: convergence analysis
Let us start with the example of the use of the two-step extrapolation procedure for the ground state of hydrogen-like xenon ($Z=54$), see Fig. 5. The many-potential term is considered, the results for both gauges are presented. At the first stage, the extrapolation with respect to the size of the basis $N$ is performed. For convenience, the extrapolation curves $k=10$, $15$ and $20$ are shown. The corresponding extrapolated sums $S_{k}$ are depicted as stars. The other extrapolated partial sums are depicted with dashes as well. At the second stage, the approximation over $k$ is performed. The final value ($\Delta\varepsilon_{a}^{mp}=S_{\infty}$, black) is obtained. Note that the GF method yields results that are indistinguishable at this scale from those obtained by the DKB method.
Based on our experience, there is no significant difference in extrapolation by basis size across two gauges. This can be seen, e.g., in the Fig. 5, where the many-potential term differ greatly (by one and a half orders of magnitude) and the rate of convergence in the basis-set size is approximately the same.
@@
Figure 5: The extrapolation of the many-potential contribution to the self-energy correction for the $1S_{1/2}$ state of hydrogen-like xenon ($Z=54$) obtained within the framework of the DKB approach in Feynman (a) and Coulomb (b) gauges.
@@
Figure 6: The many-potential contributions $\Delta\varepsilon_{a}^{\text{Mp}}(k)$: (a) normalized to the full self-energy correction $\Delta\varepsilon_{a}$ and (b) additionally multiplied by $k^{3}$. Note the log-log scale.
Let us examine the $k$-dependence of the many-potential contributions in more detail and perform the comparative analysis of the $k$ – convergence of the many-potential contribution in two gauges. The zero-, one-, and many-potential contributions to the SE correction in two gauges for $Z=18$ and $Z=92$ as a series in $k$ are presented in Tab. 1. There, at the bottom of the table, the results obtained by applying the extrapolation procedure to the corresponding PW-expansion series are shown. Note that the notation $\Delta\varepsilon_{a}^{\text{Mp}}(k)$ is introduced. The latter is the contribution to $\Delta\varepsilon_{a}^{\text{Mp}}$ corresponding to the given $k$ (thus, the partial sum $S_{k}$ for many-potential term is $S_{k}=\sum_{k^{\prime}=1}^{k}\Delta\varepsilon_{a}^{\text{Mp}}(k^{\prime})$).
@@
In Fig. 6 (a) the individual $k$ terms of the PW expansion of the many-potential contribution from Tab. 1 normalized to the full self-energy are plotted. From both the table and the graph it can be noticed that for all $Z$ both the many-potential contribution and the first PW contributions are significantly smaller in the Coulomb gauge (in the case of the $Z=18$ argon – by two orders of magnitude). This difference, however, becomes much smaller in the large-$k$ area, although the contributions are still larger in the Feynman gauge.
Multiplying by the $k^{3}$ factor allows a much closer comparison of the decay rate of contributions, see Fig. 6 (b). From our study it follows that the individual contributions are decreasing as $\Delta\varepsilon_{a}^{\text{Mp}}(k)\sim\frac{1}{k^{3}}$. That justifies that the smallest power used in $k$ extrapolation in (43) is $\frac{1}{k^{2}}$, since $\sum_{k^{\prime}=1}^{k}\frac{1}{k^{\prime 3}}\sim\frac{1}{k^{2}}$. At the same time, convergence to $\frac{1}{k^{3}}$ is clearly slower in the Coulomb gauge. An indirect interpretation is that the coefficients at the higher powers in the expansion (43) are larger in the Coulomb gauge case.
The latter circumstance apparently balances out the smallness of the many-potential contribution itself in the Coulomb gauge. In practice, given the same resource costs (number of basis functions, number of integration nodes for radial variables, etc.), Coulomb and Feynman gauges are equally suitable for estimating the many-potential term of the SE correction within a wide range of nuclear charges $Z$.
IV.2 Notes on the acceleration-convergence schemes

### v2

@@ III Computational details
The calculations are carried out by means of two approaches: the FBS method within the DKB basis and the GF method. The methods differ only in the treatment of the terms evaluated in coordinate space. The momentum-space contributions are the same in both approaches.
@@ III.1 Momentum-space calculations
The contributions evaluated in momentum space include the zero- and one-potential terms resulting from the renormalization procedure, as well as the quasi-two-potential terms corresponding to the SC convergence-acceleration scheme. All the integrals in the zero-, one-, and quasi-two-potential terms are calculated using the Gauss-Legendre quadratures. Let us first briefly discuss an evaluation of the zero-potential terms. The domain of integration over $p$ is divided into two parts, $p\leq p_{0}$ and $p\geq p_{0}$, where $p_{0}=\alpha Z$. The integration over the second domain can be carried out using a change of variables of the form $t=\alpha Z/p$, as a result of which the ray $\alpha Z,\infty)$, is mapped onto the interval $[1,0]$. The first domain is finite, so its treatment causes no problems. Compared to the Feynman gauge, the zero-potential contribution in the Coulomb gauge contains an additional integration over the Feynman parameter $x$, see Eqs. (17), (60), and (61). The coefficient $F_{2}$ in Eq. (61) must be calculated as a principal-value integral. From a numerical point of view, it is helpful to isolate the singularity. We do that by adding and subtracting the same integrand but with the numerator evaluated at the singularity point $x=a$:
@@
The first term in the right-hand side of Eq. (45) becomes regular. To avoid numerical problems at $x=a$, the exact expression can be replaced with its Taylor series in a small vicinity of this point. The second term in Eq. (45) is calculated analytically. Similar methods can be used to calculate $\Delta\tilde{\varepsilon}_{a}^{2\text{p}}$ in momentum space. To avoid possible numerical problems due to large cancellations, one should use the Taylor-series expansion for the corresponding integrands.
The evaluation of the one-potential contributions involves the integrations over the variables $p$, $p^{\prime}$, $z$, and a number of Feynman parameters. When integrating over the Feynman parameters in both gauges, to avoid possible numerical errors, the Taylor series for the integrands are employed. In the Coulomb-gauge case we use the results of Ref. Holmberg (2011), but expand all the coefficients $F_{1}$ - $F_{21}$ at ‘‘suspicious’’ points. The expression (20) has an integrable singularity for $q=0$ when $p=p^{\prime}$ and $z=1$. Therefore it is necessary to properly choose integration nodes for the variable $z$ to ensure the integral convergence. For this goal, we use the transformation $z=1-2t^{2}$ with $t\in[0,]$. To calculate the integrals over the variables $p$ and $p^{\prime}$, we change the variables according to $P=(p+p^{\prime})/\sqrt{2}$ and $P^{\prime}=(p-p^{\prime})/\sqrt{2}$. The integral over $P$ is calculated similar to the integral over $p$ for the zero-potential contribution, but the separation point $p_{0}=2\alpha Z$ is chosen instead. The integral over $P^{\prime}$ is taken within the finite interval [P,P][-P,P].
To estimate an uncertainty of the momentum-space calculations, the number of quadrature nodes and position of the separation points $p_{0}$ are varied.
@@ III.2 Coordinate-space calculations
The calculation of all coordinate-space contributions, many-potential, two-potential, and (quasi-) three-plus-potential terms, has many common points. Let us illustrate them with an example of the many-potential contribution. Its evaluation essentially reduces to the calculation of three integrals: over the radial variables $r_{1}$ and $r_{2}$, see Eqs. (75) and (76), and over the energy parameter $\omega$. Taking into account the exponential decay of bound-electron wave functions at large distances, all radial integrations are performed over the finite interval $0,R_{\text{max}}]$. The value of the parameter $R_{\text{max}}$ is chosen so that the results are independent of it within the desired accuracy. All the radial integrations are carried out using the Gauss–Legendre quadratures. However, the specific details differ slightly in the DKB and GF approaches.
In the DKB approach, the domain $0,R_{\text{max}}]$ is divided into a finite number of intervals by the nodes used to construct the B splines. The integration method for the double integral over $r_{1}$ and $r_{2}$ depends on whether these variables belong to the same interval or not. If not, the integrations are carried out independently using the same quadrature for both variables. However, when $r_{1}$ and $r_{2}$ fall into the same interval, a more delicate integration is employed to take into account the fact that the integrand depends on $r_{<}=\min\{r_{1},r_{2}\}$ and $r_{>}=\max\{r_{1},r_{2}\}$. Namely, for some fixed value of $r_{1}$ the corresponding interval is divided into two subitervals lying to the left and to the right from $r_{1}$. The integration over $r_{2}$ is performed by using the Gauss-Legendre quadratures for both subintervals. Note that within the DKB approach, the integration over radial variables turns out to be ‘‘tied’’ to the size of the basis $N$. As the FBS size grows, the number of nodes also increases, that leads to the improvement of the integration accuracy. Thus, the single parameter $N$ affects both the accuracy of the Green’s function representation and the integration accuracy.
In the GF approach, the radial integrations become more complicated due to the aforementioned discontinuities of the electron Green’s functions. The integrations over $x$ and $y$ in Eq. (28) cannot be performed independently of the integrations over $r_{1}$ and $r_{2}$. To evaluate the radial integrals, we employ a variation of the method described, e.g., in Ref. Yerokhin and Maiorova (2020). Namely, a hierarchy of the radial integration grids is prepared, with each subsequent grid becoming finer and ‘‘lying’’ inside the previous one. Then, all the radial integrations are arranged in the following order: $r_{1}$, $r_{2}$, $x$, and $y$. Therefore, the finest grid corresponds to the variable $y$. The accuracy of the GF approach is determined mainly by the accuracy of the radial integrations.
@@
Figure 4: Deformed contour (blue solid line) in the complex plane $\omega$. Singularities and branching cuts of the electron Green’s function (black lines and black dots) and photon propagator (black dashed lines) are shown.
The photon and electron propagators have a complicated analytical structure as functions of the energy parameter $\omega$ in the complex plane. They are two-valued functions and are defined on the corresponding Riemann surfaces with two sheets (for each propagator). Fig. 4 shows the complex plane of the integration variable $\omega$, the branch points of both propagators with the cuts starting from them, and the poles of the electron propagator corresponding to the bound states. In the expression for the SE correction (3), the integration contour goes along the real axis, but it can be bended in the $\omega$ complex plane to ‘‘simplify’’ the calculations. Because of the zero photon mass, the contour is squeezed between two infinitely close branching points near zero. Introducing a finite mass would allow them to be pushed apart, which can be easily demonstrated in the Feynman gauge. Consequently, the contour must always pass through $\omega=0$. In principle, there is no other restrictions on the contour. Thus, it can be changed as needed, respecting the analytical structure of the integrand.
In this work, the approach proposed in Ref. Yerokhin and Shabaev (1999) is used. Namely, we employ the transformation that results in the contour shown in Fig. 4. The original contour along the real axis is rotated and deformed. The contributions from the circular arcs vanish as the radius of the circle tends to infinity, so we do not consider them. The rotation is done to avoid fast oscillations of the integrand encountered on the real axis: on the vertical parts of the deformed contour the integrand decays exponentially. The deformation of the contour near zero and its shift to the right from the imaginary axis is done to avoid going close to the poles of the electron Green’s function, which correspond to the bound states with binding energies no less than that of the state under consideration. For large enough values of $\kappa$, the contour simplifies and can be taken along the imaginary axis. The integration over $\omega$ is evaluated using the Simpson quadratures. The number of integration points is increased until the required accuracy is achieved.
@@ III.3 Extrapolation of the PW-expansion sums
We obtain a coordinate-space contribution as a PW series in $\kappa$, which should inevitably be truncated at some $|\kappa_{\text{max}}|$. The remainder of the series can then be estimated using an extrapolation procedure. The extrapolation methods discussed below are applied to all treated coordinate-space contributions. Here and in what follows, when discussing partial sums, we will denote by the index $k>0$ the sum of all terms of the PW expansion with $|\kappa|\leq k$.
When using the DKB method, the individual contributions strongly depend on the FBS size $N$, therefore, an extrapolation over this parameter is also required. In this case, the most stable results are obtained if one extrapolates with respect to the basis size first. Let us denote as $S_{k}(N)$ the partial sum obtained for the given FBS size $N$. For each $k$, the extrapolation over $N$ is performed using a non-linear least-squares method using an ansatz of the form:
@@
where $S_{k}$ are the extrapolated to the infinite basis set partial sums, and $A_{k}$ and $B_{k}$ are adjustable numerical coefficients. In the GF approach, the sums $S_{k}$ are obtained directly from the calculations, and this extrapolation is not required.
To extrapolate over $k$ in both DKB and GF approaches, we implement three following ansatzes for the partial sums $S_{k}$:
@@
where the parameter $S_{\infty}$ and coefficients $C_{i}$ for $i=2,\ldots,m$ (with $m$ being $4$, 55 and $6$) are obtained by a kind of statistical method. Namely, for each $m$, $N_{\text{iter}}$ times a subset of $k$’s, $\{k_{i}\}_{i=1}^{m}$, is randomly selected. The parameter $N_{\text{iter}}$ is fixed by the condition that all possible subsets are considered for the largest $m$ ($m=6$). For each element of the above subsets we construct and solve the system of $m$ linear equations of the form
$\tilde{S}(k_{i})=S_{k_{i}},\quad i=1,\dots,m$ (48)
to determine $m$ variables $S_{\infty}$, $C_{2}$, …, $C_{m}$. Repeating this step $3N_{\text{iter}}$ times yields a set of approximations for $S_{\infty}$. Evaluating the mean value $\overline{S}_{\infty}$ and variance $\sigma_{k}$, one obtains the desired extrapolated sum and an estimation for its uncertainty. We note that we approximate the tails only (by excluding the first few $k$’s from consideration). Changing the tail size produces results that are consistent with each other.
Within the DKB approach, the uncertainties of extrapolation in $k$, $\sigma_{k}$, and extrapolation to an infinite basis set, $\sigma_{N}$, are summed quadratically
@@
Example of the use of this two-step extrapolation procedure can be found in the ‘‘results’’ section below.
@@ IV Results
In the present work, we conduct two sets of SE-correction calculations. First, we consider the $1S_{1/2}$ state in three hydrogen-like ions: argon ($Z=18$), xenon ($Z=54$), and uranium ($Z=92$). We adopt the results from Ref. Hedendahl and Holmberg (2012) as a benchmark for comparison. For this reason, in the first part of calculations, the values of root-mean-square (rms) nuclear radii as well as the values of fundamental constants, e.g., the fine-structure constant $\alpha$ and the Hartree energy in eV, are taken to be the same as there. The obtained data are used to analyze the convergence of the many-potential contribution. Second, we study the SE corrections for the $1S_{1/2}$, $2S_{1/2}$, $2P_{1/2}$, and $2P_{3/2}$ states in a set of hydrogen-like ions with the nuclear charge $Z$ in the range from $Z=10$ to $Z=100$ and compare the obtained results with the ones presented in Ref. Yerokhin and Shabaev (2015a). In this case, the rms radii and models of the nuclear-charge distribution are chosen to coincide with those in the corresponding work, however, for $\alpha$ the numerical value from the CODATA (2018) Tiesinga et al. (2021) is used. In the second part, the SE correction is represented in terms of the dimensionless function $F(\alpha Z)$, defined as:
@@
The data obtained in the second part of the work are used to analyze the acceleration-convergence schemes.
@@ IV.1 Many-potential term: convergence analysis
Let us start with illustrating the two-step extrapolation procedure, described in the previous section, using the ground state of hydrogen-like xenon ($Z=54$) as an example, see Fig. 5. The many-potential term is considered, and the results for both gauges are presented. At the first stage, the extrapolation of the partial sums $S_{k}(N)$ with respect to the size of the basis $N$ is performed. For convenience, only the extrapolation curves for $k=10$, $15$, and $20$ are shown as functions of $1/N$. The corresponding extrapolated sums $S_{10}$, $S_{15}$, and $S_{20}$ are depicted as stars at $1/N\rightarrow 0$. The extrapolated partial sums $S_{k}$ obtained for other values of $k$ are depicted with dashes as well. At the second stage, the extrapolation over $k$ is performed. The final values for $\Delta\varepsilon_{a}^{\text{Mp}}=S_{\infty}$ are shown with black squares. Note that the GF method yields results that are indistinguishable at this scale from those obtained by the DKB method. Based on our experience, a small difference in the FBS size extrapolation in the two gauges appears in the small-$Z$ region, where the convergence behavior is slightly better in the Coulomb gauge. In the region of medium- and large-$Z$ there is no significant difference between the gauges. This can be seen, e.g., from Fig. 5, where the absolute values of the many-potential terms in two gauges differ by approximately one order of magnitude, while the rate of convergence in the FBS size is approximately the same.
@@
Figure 5: The extrapolation of the many-potential contribution to the self-energy correction for the $1S_{1/2}$ state of hydrogen-like xenon ($Z=54$) obtained within the framework of the DKB approach in the Feynman (a) and Coulomb (b) gauges.
@@
Figure 6: The individual many-potential contributions $\Delta\varepsilon_{a}^{\text{Mp}}(k)$: (a) normalized to the total self-energy correction $\Delta\varepsilon_{a}$ and (b) additionally multiplied by $k^{3}$. Note the log-log scale.
Let us now examine in details the PW convergence of the many-potential contributions and perform a comparative analysis within two gauges. The zero-, one-, and many-potential contributions to the SE correction for the $1S_{1/2}$ state in hydrogen-like argon ($Z=18$) and uranium ($Z=92$) in Feynman and Coulomb gauges are presented in Table 1. For the many-potential contributions, the individual terms of the PW expansion, $\Delta\varepsilon_{a}^{\text{Mp}}(k)$, obtained by means of the GF approach are shown. These terms represent sums of contributions with $\kappa=k$ and $\kappa=-k$, i.e., the partial sums are given by $S_{k}=\sum_{i=1}^{k}\Delta\varepsilon_{a}^{\text{Mp}}(i)$. The row labeled $\sum_{k=1}^{24}$’’ gives the partial sums $S_{24}$, while the subsequent row $\sum_{k=25}^{\infty}$ [extr.]’’ provides our estimations for the remainders of PW series obtained by applying the statistical extrapolation procedure discussed above. The total SE corrections are presented in the last line.
@@
For convenience, the absolute values of the individual terms $\Delta\varepsilon_{a}^{\text{Mp}}(k)$ from Table 1, normalized to the total self-energy correction $\Delta\varepsilon_{a}$, are plotted in Fig. 6(a). From Table 1 and Fig. 6(a), it can be noticed that for both considered $Z$ the many-potential contributions as well as the first terms of the PW expansions are significantly smaller in the Coulomb gauge (in the case of $Z=18$, by two orders of magnitude) than in the Feynman gauge. This difference, however, becomes much less noticeable for large values of $k$, although the Feynman-gauge contributions still exceed their Coulomb-gauge counterparts.
Multiplying the individual terms $\Delta\varepsilon_{a}^{\text{Mp}}(k)$ obtained for two gauges by the factor of $k^{3}$ provides a more informative comparison of their decay rate with the growth of $k$, that is demonstrated in Fig. 6(b). Let us discuss the leading asymptotic behavior of the PW-expansion terms for $k\rightarrow\infty$: $\Delta\varepsilon_{a}^{\text{Mp}}(k)\sim 1/{k^{p}}$. Our numerical results indicate that in both the Feynman and Coulomb gauges, the exponent $p$ exceeds 2 over a wide range of $Z$, with $p$ being close to $3$ in the large-$Z$ region. These conclusions do not contradict those from Ref. Yerokhin et al. (2005), where it was claimed that in the Feynman gauge $p=3$ regardless of the nuclear charge $Z$. The latter observation justifies the fact that the anzatz for the extrapolation over $k$ in (47) starts from ${1}/{k^{2}}$, since $\sum_{i=1}^{k}{1}/{i^{3}}\sim{1}/{k^{2}}$.
We also note that the numerical coefficients for the higher powers of $1/k$ in the expansion (47) in the Coulomb gauge are larger than their Feynman-gauge counterparts. That fact is also illustrated in the Fig. 6(b) – convergence to asymptotics in the Feynman gauge is significantly slower. As a result, within the $k$-range studied in the present work ($k\leq 20$ or $k\leq 24$), the leading asymptotic behavior is less pronounced in the Coulomb gauge. The latter fact apparently balances out the smallness of the absolute value of the Coulomb-gauge many-potential contribution. In practice, within the GF approach, given the same resource costs (number of partial waves, number of integration nodes for radial variables, etc.), the Coulomb and Feynman gauges are equally suitable for estimating the many-potential contribution to the SE correction within a wide range of nuclear charges $Z$. This conclusion slightly contradicts what was claimed in Ref. Yerokhin et al. (2025a). Nevertheless, we agree with the fact that the advantages of the Coulomb gauge are the improved convergence of the many-potential contributions over the basis-set size in the small-$Z$ region, relative smallness of the many-potential contribution Hedendahl and Holmberg (2012); Yerokhin et al. (2025a), and the absence of cancellation between different parts of the SE correction.
IV.2 Application of the convergence-acceleration schemes